OpenStudy (anonymous):

find (hog)(1) for g(x)=-3x, h(x)=x^3+2

OpenStudy (anonymous):

hog(x) =h(g(x) = h(-3x) =(-3x)^3 + 2 = -27x^3 + 2 then hog(1) = - 27(1)^3 + 2 = - 27 + 2 = - 25